Sun 31 Aug - Power Outage
All power to the campus was switched-off for a period last Sunday (31st August), to allow for essential maintenance work to be carried out.
The Interactive UNIX Service (IUS) was unavailable during this time.
On the successful completion of the work, power was restored. No related IT problems were reported.
Fri 29 Aug - Office 2007 Deployment to PC Labs
ISS commenced deployment of Office 2007 to ISS bookable Labs, Lecture Theatres and Seminar Rooms.
Thu 28 Aug - All systems running without problems
Wed 27 Aug - Exchange Email Service Maintenance
The Exchange email system was temporarily unavailable whilst security updates were applied.
Tue 26 August - All systems ran without problems
Mon 25 August - University closed
Fri 22 Aug - All systems ran without problems
Thu 21 Aug - Office 2007 Deployment to PC Labs
ISS deployed Office 2007 ISS managed PC Labs on Thursday 21st, August. This deployment included:
* ISS non-bookable Labs listed at http://www.lancs.ac.uk/iss/pclabs/
* ISS open access PCs in the Library
* The PCs in the Library's Training Suites
* Open access PCs in college foyers and their PC Labs
* PC Labs and Base Rooms in the Management School.
Wed 20 Aug - LUVLE Services Maintenance & Biology and Environmental Network at Risk
LUVLE Services Maintenance
LUVLE, myPlace, my Modules, online assessments and module feedback were unavailable for an extended downtime to allow for the replacement of server hardware and essential server maintenance.
Biology and Environmental Network at Risk
Due to refurbishment works in Biology a short network outage occurred at around 11:00 this morning while the equipment was moved onto a temporary power feed. Further work was necessary to move the equipment back onto a proper power feed which caused a 3 to 5 minute outage.
Tue 19 Aug - All systems ran without problems
Mon 18 Aug - All systems ran without problems
Fri 15 Aug - All systems ran without problems
Thu 14 Aug - All systems ran without problems
Wed 13 Aug - All systems ran without problems
Tue 12 Aug - All systems ran without problems
Mon 11 Aug - All systems ran without problems
Fri 08 Aug - HPC head node rebooted
The HPC corona head node was rebooted to cure a problem with access to files in home directories, which was affecting some users. Submitted jobs continued to run.
Thu 07 Aug - Network Maintenance - University House
There was a slight disruption to the network in University House to allow for installation of new network equipment for the refurbishment of the B floor. Downtime was kept to a minimum and lasted no more than 5 minutes while this critical work took place.